The past five monthly returns for PG&E are −3.47 percent, 4.63 percent, 4.07 percent, 6.92 percent, and 3.88 percent. Compute the standard deviation of PG&E’s monthly returns. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Average return=Total return/Total time period

=(-3.47+4.63+4.07+6.92+3.88)/5=3.206%

Return (Return-Average Return)^2
-3.47 (-3.47-3.206)^2=44.568976
4.63 (4.63-3.206)^2=2.027776
4.07 (4.07-3.206)^2=0.746496
6.92 (6.92-3.206)^2=13.793796
3.88 (3.88-3.206)^2=0.454276
Total=61.59132%

Standard deviation=[Total (Return-Average Return)^2/(Time period-1)]^(1/2)

=[61.59132/(5-1)]^(1/2)

=[61.59132/4]^(1/2)

=3.92%(Approx).